OpenStudy (anonymous):

Describe how the discriminant can help you determine the type of solutions you are going to have NEED HELP FAST!!!!!

OpenStudy (unklerhaukus):

A Quadratic Equation\[ax^2+bx+c=0\] The Quadratic Formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] The Discriminant is \(\Delta =b^2-4ac\) if \(\Delta=0\) there is only one solution to the quadratic equation if \(\Delta>0\) there are two solutions if \(\Delta <0\) the solutions are complex

OpenStudy (unklerhaukus):

because the square root of zero is zero, there are two square roots of a positive number, and the two complex square roots to a negative number